# Quantum reference frames for spacetime symmetries and large gauge transformations

### v1

@@ 2 Operational QRFs
$(\mathcal{M}_{\mathcal{S}}\otimes B(\mathcal{H}_{\mathcal{R}}))^{G}=(\operatorname{id}_{\mathcal{H}_{\mathcal{S}}}\otimes V)^{*}((\mathcal{M}_{\mathcal{S}}\otimes L^{2}(G))^{G}\otimes B(\mathcal{K}))(\operatorname{id}_{\mathcal{H}_{\mathcal{S}}}\otimes V),$ (2)
@@
This is a special case of [11, Thm. 4.9]. Here the space $\mathcal{K}$ and map $V$ can be constructed explicitly from the QRF using a group representation theoretic result known as Mackey’s imprimitivity theorem [17, 18]. The algebra $(\mathcal{M}_{\mathcal{S}}\otimes L^{2}(G))^{G}$ is a well-understood object in the theory of von Neumann algebras and goes under the name crossed product algebra $\mathcal{M}_{\mathcal{S}}\rtimes_{\alpha}G$ [19]. Crossed product algebras have become a recent focus in theoretical physics due to their appearance in models for quantum gravity, where they play a role in defining a notion of entropy [20, 21]. Below, we describe a generalisation of these results in the context of QFT on curved spacetimes combined with QRFs.
@@ 4 Gauge theories, boundaries and quantum reference frames
At each Cauchy surface with boundaries $\overline{\Sigma}$, the observables of the theory can be expressed in terms of smearings of the EM vector potential $\mathbf{A}$ and electric field $\mathbf{E}$ (understood as differential one-forms on $\overline{\Sigma}$). These smearings are such that the observables are invariant under gauge transformations $\mathbf{A}\mapsto\mathbf{A}+\textup{d}\Lambda$, where $\Lambda$ is a smooth function on $\overline{\Sigma}$ vanishing on its boundary $\partial\overline{\Sigma}$. This notion of gauge invariance is sufficient to ensure well-behaved dynamics of the theory. The resulting algebra contains observables that transform covariantly under boundary/large gauge transformations (i.e. for which $\Lambda$ does not vanish on $\partial\overline{\Sigma}$). These are dubbed edge-mode observables and can be understood as (smeared analogues of) Wilson lines connecting two points on the boundary $\int_{0}^{1}\mathbf{A}_{\mu}\textup{d}\gamma^{\mu}$ for a curve $\gamma:[0,1]\to\overline{\Sigma}$ with $\gamma(0),\gamma(1)\in\partial\overline{\Sigma}$. By quantising the Poisson structure arising naturally from the covariant phase space formalism for electromagnetism on spacetimes with boundaries [34], these edge mode observables have non-trivial commutation relations with smearings of the electric flux density $\textup{n}_{\partial\overline{\Sigma}}\mathbf{E}$ through $\partial\overline{\Sigma}$. As a consequence, these fluxes are quantised, rather than superselected as in conventional approaches to quantum electromagnetism [35].

### v2

@@ 2 Operational QRFs
$(\mathcal{M}_{\mathcal{S}}\otimes B(\mathcal{H}_{\mathcal{R}}))^{G}=(\operatorname{id}_{\mathcal{H}_{\mathcal{S}}}\otimes V)^{*}((\mathcal{M}_{\mathcal{S}}\otimes B(L^{2}(G)))^{G}\otimes B(\mathcal{K}))(\operatorname{id}_{\mathcal{H}_{\mathcal{S}}}\otimes V),$ (2)
@@
This is a special case of [11, Thm. 4.9]. Here the space $\mathcal{K}$ and map $V$ can be constructed explicitly from the QRF using a group representation theoretic result known as Mackey’s imprimitivity theorem [17, 18]. The algebra $(\mathcal{M}_{\mathcal{S}}\otimes B(L^{2}(G)))^{G}$ is a well-understood object in the theory of von Neumann algebras and goes under the name crossed product algebra $\mathcal{M}_{\mathcal{S}}\rtimes_{\alpha}G$ [19]. Crossed product algebras have become a recent focus in theoretical physics due to their appearance in models for quantum gravity, where they play a role in defining a notion of entropy [20, 21]. Below, we describe a generalisation of these results in the context of QFT on curved spacetimes combined with QRFs.
@@ 4 Gauge theories, boundaries and quantum reference frames
At each Cauchy surface with boundaries $\overline{\Sigma}$, the observables of the theory can be expressed in terms of smearings of the EM vector potential $\mathbf{A}$ and electric field $\mathbf{E}$ (understood as differential one-forms on $\overline{\Sigma}$). These smearings are such that the observables are invariant under gauge transformations $\mathbf{A}\mapsto\mathbf{A}+\textup{d}\Lambda$, where $\Lambda$ is a smooth function on $\overline{\Sigma}$ vanishing on its boundary $\partial\overline{\Sigma}$.
@@
This notion of gauge invariance is sufficient to ensure well-behaved dynamics of the theory. The resulting algebra contains observables that transform covariantly under boundary/large gauge transformations (i.e. for which $\Lambda$ does not vanish on $\partial\overline{\Sigma}$). These are dubbed edge-mode observables and can be understood as (smeared analogues of) Wilson lines connecting two points on the boundary $\int_{0}^{1}\mathbf{A}_{\mu}\textup{d}\gamma^{\mu}$ for a curve $\gamma:[0,1]\to\overline{\Sigma}$ with $\gamma(0),\gamma(1)\in\partial\overline{\Sigma}$. By quantising the Poisson structure arising naturally from the covariant phase space formalism for electromagnetism on spacetimes with boundaries [34], these edge mode observables have non-trivial commutation relations with smearings of the electric flux density $\textup{n}_{\partial\overline{\Sigma}}\mathbf{E}$ through $\partial\overline{\Sigma}$. As a consequence, these fluxes are quantised, rather than superselected as in conventional approaches to quantum electromagnetism [35].